# A Robust Real-Time Lane Detection Method with Fog-Enhanced Feature Fusion for Foggy Conditions

### v10

@@ 4.1 Overall structure
To address the challenge of lane detection in foggy conditions, where visibility is significantly reduced, we propose a fog-enhanced lane detection network. The network works through several collaborative modules to tackle the challenges posed by foggy images, with the overall architecture shown in Fig. 8. The input image first passes through the backbone network, where feature extraction occurs, followed by the Global Feature Fusion Module (GFFM). This module is designed to mitigate the loss of global information in foggy images, allowing the network to retain essential global context. The output is a feature map enriched with both local and global information, crucial for accurate lane detection.
The multi-level feature map from the backbone is then passed into the Neck section of the network, which employs HRNetV2[47] for multi-scale feature fusion. HRNetV2 excels in integrating high-resolution and low-resolution features from different stages of the backbone. This process helps the network capture spatial information at multiple scales, producing a rich, multi-scale feature representation that forms the basis for lane instance detection.
@@
The output parameter map is then passed to the Kernel Feature Fusion Module (KFFM), which learns the relationships between different lane instances. The KFFM fuses these instance-specific features, improving lane localization accuracy, especially for complex or overlapping lane lines. This module ensures that the network can better identify lane instances and refine the localization of each lane.
Finally, to compensate for the loss of edge information in foggy images, the network incorporates a Low-Level Edge Enhancement Module (LEEM). This module enhances edge features using convolution operations, recovering fine details at lane boundaries. The enhanced edge features improve the network’s ability to detect lane boundaries with high precision, even under challenging foggy conditions.
Figure 8: Overview of the proposed network.
@@ 4.2 Global Feature Fusion Module
As shown in Fig. 9, to enhance global feature integration in the Swin Transformer backbone[48], we replace the final window-based attention mechanism with a standard multi-head attention mechanism (MHSA) from Vision Transformer (ViT)[49]. While ViT excels at capturing global dependencies, it suffers from high computational cost, especially at high resolutions, making it unsuitable for real-time tasks such as lane detection. The traditional ViT computes global self-attention with a complexity of $O(N^{2})$, where $N$ is the number of input tokens. This high computational cost becomes prohibitive at high resolution, slowing down inference and limiting real-time processing.
In our approach, we address this issue by introducing an efficient solution: global attention is applied only after reducing the image resolution. This ensures that the computational cost of global attention is manageable, allowing us to retain the global attention mechanism of ViT while maintaining fast inference speed, which is crucial for real-time lane detection in foggy conditions.
@@
Figure 9: Backbone Network with Global Feature Fusion.
4.3 Kernel Feature Fusion Module
Existing lane instance segmentation methods[50, 36], typically use dynamic convolution to generate kernel parameters for each instance, which improves the network’s flexibility and adaptability. However, these methods do not fully leverage the structural relationships between lane instances, leading to limited segmentation and detection accuracy. Traditional dynamic convolution methods only consider individual lane instances and overlook the potential relationships between adjacent instances, which reduces their performance when handling complex lane scenes.
To address this limitation, we propose the Kernel Feature Fusion Module (KFFM), as shown in the bottom right of Fig. 8, which optimizes kernel generation by integrating information from adjacent lane instances. Specifically, KFFM introduces structured feature fusion into dynamic convolution, allowing each lane instance’s kernel to depend not only on the features of the current lane but also on the information from nearby lane instances. This approach enables the network to adaptively learn the relationships between lane instances, improving both detection accuracy and segmentation performance.
@@ 4.3 Kernel Feature Fusion Module
4.4 Low-level Edge Enhanced Module
Images captured in foggy conditions differ from those captured in good lighting primarily due to blurriness and low contrast, which makes the edges of objects and scene details indistinct and difficult to detect. This degradation of edge information significantly impacts the algorithm’s ability to accurately understand and recognize the scene. Edge information, beyond simply marking object boundaries, also implicitly contains structural and spatial relationships about objects, which is critical for tasks such as object detection, segmentation, and tracking. We used the Canny edge detection algorithm with identical threshold settings to process both foggy and normal images, as shown in Fig. 10. The results show that fog significantly degrades edge information, particularly lane line edges, which become nearly indiscernible. Given that lane detection relies heavily on edge information—lane positioning is typically determined by the boundary between lane edges and the road background—accurate edge detection is essential for reliable lane detection.
To address the challenges posed by foggy conditions, we propose the integration of global feature fusion within both the backbone network and detection head to enhance high-dimensional global features. Additionally, we incorporate multi-scale fusion to aggregate detailed features from low-resolution feature maps. However, this approach alone does not fully compensate for the loss of edge details in foggy images. To resolve this issue, we introduce the Low-level Edge Enhanced Module (LEEM) into the network. LEEM focuses on edge information, effectively recovering and refining the edges that are crucial for lane detection, thus improving both the accuracy and stability of lane detection in foggy environments.
@@ 4.4 Low-level Edge Enhanced Module
LEEM works by adding an auxiliary path to the network that specifically learns edge features of the input image. The process begins with the application of the Canny operator for edge detection, followed by the reconstruction of missing lane edges using lane line labels. This ground truth for the edge enhancement module is generated by combining detected edges with labeled edges. This approach helps the network better identify lane boundaries, which are often difficult to distinguish in foggy conditions.
@@
$\hat{\mathit{E}}_{\mathit{predict}}=\mathit{Conv}_{2}\left(\mathit{ReLU}\left(\mathit{BN}(\mathit{Conv}_{1}(\mathit{X})))\right)\right)$ (6)
@@
The LEEM module allows the network to effectively enhance critical edge details in lane detection, significantly improving both the accuracy and stability of lane detection in foggy conditions. This module not only recovers lost edge information in images but also integrates information from various sources, using the cross-entropy loss function to optimize the network. As a result, LEEM plays a key role in enhancing the robustness of lane detection systems in challenging environments.
[…]
  • 1.
Figure 11: Visualization results of CLRNet[16], FLAMNet[32], PolarRCNN[51] and our method on FoggyLane testing set. The arrows indicate instances of lane lines that were not correctly detected by the corresponding method.
[…]
  • 1.
@@ 5.1 Datasets
In the Data Construction chapter, we introduce the FoggyLane dataset, which was constructed from real-world collections for foggy lane detection, along with the FoggyCULane and FoggyTusimple datasets that were synthesized through modeling foggy images. Our method is evaluated on these three datasets.
FoggyLane is a lane detection dataset specifically designed for foggy scenarios, containing 1,423 frames of foggy lane images for training and testing. All images are of resolution 1640x590 pixels.
FoggyCULane is synthesized from the large-scale challenging lane detection dataset CULane using a foggy image modeling approach. It maintains the same number of images as CULane. The resolution of all images is also 1640x590 pixels.
FoggyTusimple is synthesized from Tusimple through the foggy image modeling method. At the same time, the dataset follows the same structure as Tusimple, comprising 3,268 training images, 358 validation samples, and 2,782 test cases. All visual data maintains a consistent resolution of 1280×720 pixels.
@@ 5.2 Evaluation Metric
Lane detection is typically approached as a semantic segmentation problem, assessed using common metrics. However, due to its distinct characteristics, lane detection also requires specialized evaluation metrics. Adhering to SCNN’s [8] framework, we link labeled or predicted lane points to form line segments and extend them into 30-pixel-wide regions. Subsequently, we compute the Intersection over Union (IoU), and if the IoU value exceeds a predefined threshold, typically set at 0.5, the prediction is considered correct. From this, we calculate Precision, Recall, and F1-score using the following formulas:
@@
Figure 12: Visualization results of CLRNet[16], FLAMNet[32], PolarRCNN[51] and our method on FoggyCULane testing set. The arrows indicate instances of lane lines that were not correctly detected by the corresponding method.
@@ 5.3 Implementation Details
To benchmark against existing methods[8, 35, 36, 25, 31, 28, 16, 32, 29, 51], we followed the training settings provided in the respective papers. For FoggyCULane and FoggyTusimple, we adopted the training parameters from the original CULane and Tusimple datasets, while for FoggyLane, we extended the training epochs to ensure reliable results
@@ 5.5 Ablation Study
We conducted ablation studies on the FoggyLane and FoggyCULane datasets to evaluate the effectiveness of the GFFM, KFFM, and LEEM modules, using Swin-t as the backbone. As shown in Tables 5 and 6, each module improves the model’s performance. The GFFM module contributes the most, increasing the F1-score by 1.07 on FoggyLane and 0.41 on FoggyCULane, highlighting its ability to capture both local and global information. Adding the KFFM module further boosts the F1-score by 0.63 on FoggyLane and 0.26 on FoggyCULane. Finally, the LEEM module improves edge detection, increasing the F1-score by 0.28 on FoggyLane and 0.14 on FoggyCULane.
@@
  Model GFFM KFFM LEEM F1-score Gflops(G)
@@ 5.6 Model deployment
  Model GFFM KFFM LEEM F1-score Gflops(G)
@@
Figure 14: A edge computing system built with NVIDIA Jetson AGX Orin. After experimental validation, it has been demonstrated that our method can run accurately on the vehicle’s edge computing platform with limited computational resources, achieving a performance of 38.4 FPS.
@@ 6 Conclusion
This study focuses on lane detection in foggy conditions, aiming to address the range of challenges posed by foggy environments, which is crucial for ensuring driving safety in fog and enhancing Advanced Driver Assistance System (ADAS) capabilities. To address the current lack of publicly available datasets for lane detection in foggy conditions, we constructed a real-world foggy lane dataset, FoggyLane, and synthesized two additional foggy datasets, FoggyCULane and FoggyTusimple, based on the popular CULane and Tusimple datasets.
In response to the challenges of lane detection in foggy scenarios, we propose an end-to-end fog-enhanced lane detection method. First, this method incorporates a Global Feature Fusion Module in the backbone network to address the absence of global spatial information in foggy scenes. Second, we employ a dynamic convolution-based detection head for instance-sensitive lane segmentation, integrating a Kernel Feature Fusion Module within the detection head to automatically capture associations between lane instances. Additionally, a Low-Level Edge Enhancement Module was added to increase the network’s focus on edge information, mitigating the loss of edge details and improving lane detection accuracy under foggy conditions.
Comparative experiments on the FoggyLane, FoggyCULane, and FoggyTusimple datasets with other advanced lane detection methods demonstrate that our approach achieves superior performance in both clear and foggy conditions, significantly improving lane detection accuracy. Furthermore, with TensorRT acceleration, our method achieves an inference speed of 38.4 FPS on the NVIDIA Jetson AGX Orin, further validating its practical applicability in real-world scenarios.

### v11

@@ 3.1 Overall structure
To address the challenge of lane detection in foggy conditions, where visibility is significantly reduced, we propose PDT-Net, a robust Prior-Guided Dynamic Feature Fusion Transformer framework. The network works through several collaborative modules to tackle the challenges posed by foggy images, with the overall architecture shown in Fig. 2. The input image first passes through the backbone network, where feature extraction occurs, followed by the Global Feature Fusion Module (GFFM). This module is designed to mitigate the loss of global information in foggy images, allowing the network to retain essential global context. The output is a feature map enriched with both local and global information, crucial for accurate lane detection.
The multi-level feature map from the backbone is then passed into the Neck section of the network, which employs HRNetV2(9052469) for multi-scale feature fusion. HRNetV2 excels in integrating high-resolution and low-resolution features from different stages of the backbone. This process helps the network capture spatial information at multiple scales, producing a rich, multi-scale feature representation that forms the basis for lane instance detection.
@@
The output parameter map is then passed to the Dynamic Feature Fusion Module (DFFM), which learns the relationships between different lane instances. The DFFM fuses these instance-specific features, improving lane localization accuracy, especially for complex or overlapping lane lines. This module ensures that the network can better identify lane instances and refine the localization of each lane.
Finally, to compensate for the loss of edge information in foggy images, the network incorporates a Prior-Guided Edge Enhancement Module (PEM). This module enhances edge features using convolution operations, recovering fine details at lane boundaries. The enhanced edge features improve the network’s ability to detect lane boundaries with high precision, even under challenging foggy conditions.
@@ 3.2 Global Feature Fusion Module
As shown in Fig. 3, to enhance global feature integration in the Swin Transformer backbone(liu2021swin), we replace the final window-based attention mechanism with a standard multi-head attention mechanism (MHSA) from Vision Transformer (ViT)(vit). While ViT excels at capturing global dependencies, it suffers from high computational cost, especially at high resolutions, making it unsuitable for real-time tasks such as lane detection. The traditional ViT computes global self-attention with a complexity of $O(N^{2})$, where $N$ is the number of input tokens. This high computational cost becomes prohibitive at high resolution, slowing down inference and limiting real-time processing.
In our approach, we address this issue by introducing an efficient solution: global attention is applied only after reducing the image resolution. This ensures that the computational cost of global attention is manageable(AAP2), allowing us to retain the global attention mechanism of ViT while maintaining fast inference speed, which is crucial for real-time lane detection in foggy conditions.
Figure 4: The structure of the Dynamic Feature Fusion Module(DFFM)
@@
3.3 Dynamic Feature Fusion Module
Existing lane instance segmentation methods(tian2020conditional; liu2021condlanenet), typically use dynamic convolution to generate kernel parameters for each instance, which improves the network’s flexibility and adaptability. However, these methods do not fully leverage the structural relationships between lane instances, leading to limited segmentation and detection accuracy. Traditional dynamic convolution methods only consider individual lane instances and overlook the potential relationships between adjacent instances, which reduces their performance when handling complex lane scenes. To address this limitation, we propose the Dynamic Feature Fusion Module (DFFM), as shown in Fig. 4, which optimizes kernel generation by integrating information from adjacent lane instances. Specifically, DFFM introduces structured feature fusion into dynamic convolution, allowing each lane instance’s kernel to depend not only on the features of the current lane but also on the information from nearby lane instances. This approach enables the network to adaptively learn the relationships between lane instances, improving both detection accuracy and segmentation performance.
@@ 3.3 Dynamic Feature Fusion Module
3.4 Prior-Guided Edge Enhancement Module
In foggy conditions, images are typically characterized by reduced contrast and sharpness, making it difficult to detect key features such as edges5, which are critical for tasks like lane detection. In these scenarios, edge information becomes an essential prior knowledge that can guide the network in recovering vital details. Edge information not only defines object boundaries but also conveys structural and spatial relationships, crucial for understanding the scene. Recognizing this, we propose the Prior-Guided Edge Enhancement Module (PEM),as shown in Fig. 6, which integrates prior knowledge of edge structures to recover lost edge information in foggy images, ultimately enhancing lane detection accuracy and robustness.
@@ 3.4 Prior-Guided Edge Enhancement Module
PEM works by adding an auxiliary path to the network that specifically learns edge features of the input image. The process begins with the application of the Canny operator for edge detection, followed by the reconstruction of missing lane edges using lane line labels. This ground truth for the edge enhancement module is generated by combining detected edges with labeled edges. This approach helps the network better identify lane boundaries, which are often difficult to distinguish in foggy conditions.
@@
$\hat{\mathit{E}}_{\mathit{predict}}=\mathit{Conv}_{2}\left(\mathit{ReLU}\left(\mathit{BN}(\mathit{Conv}_{1}(\mathit{X}))\right)\right)$ (3)
@@
The PEM module allows the network to effectively enhance critical edge details in lane detection, significantly improving both the accuracy and stability of lane detection in foggy conditions. This module not only recovers lost edge information in images but also integrates information from various sources, using the cross-entropy loss function to optimize the network. As a result, PEM plays a key role in enhancing the robustness of lane detection systems in challenging environments.
Figure 6: The structure of the Prior-Guided Edge Enhancement Module(PEM)
@@ 3.5.5 Edge prediction
Figure 7: The route for dataset acquisition in Suzhou on the morning of October 26, 2023.(googlemap)
Figure 8: The routes for dataset acquisition in Urumqi, Xinjiang Autonomous Region: the left illustrates the morning of December 9, 2023, while the right depicts the evening of January 2, 2024.(googlemap)
@@ 4.1 Datasets
Currently, there is no publicly available foggy weather lane detection dataset. To fill this gap, we created the FoggyLane dataset for lane detection in foggy weather. In addition, we generated two supplementary datasets, FoggyCULane and FoggyTusimple, by using fog modeling techniques on existing CULane (pan2018spatial) and Tusimple (Tusimple) datasets to test the generalizability of our methods.
Figure 9: The different road types of FoggyLane images examples.
Dataset Frames Multi-traffic scenarios Instance-level annatation Number of foggy datasets Resolution Caltech-Lanes (Caltech-Lanes) 1.2k None 640 x 480 Tusimple (Tusimple) 6.4k None 1280 x 720 LLAMAS (LLAMAS) 100k None 1276 x 717 Appolloscape (Apolloscape) 110k None 3384 x 2710 BDD100k (yu2020bdd100k) 100k None 1280 x 720 CULane (pan2018spatial) 133k None 1640 x 590 CurveLanes (xu2020curvelane) 150k None 1280 x 720 FoggyLane (ours) 1.423k 1.423k 1640 x 590
4.1.1 FoggyLane
Existing lane detection algorithms for intelligent connected vehicle ADAS systems perform well under favorable weather conditions, but their performance significantly deteriorates in adverse conditions like fog, which reduces visibility and increases the likelihood of traffic accidents. Current benchmark datasets, including CULane(pan2018spatial), Tusimple(Tusimple), and LLAMAS(LLAMAS), lack foggy weather data. To address this, we developed FoggyLane, a dataset specifically designed for foggy conditions. Data was collected using a monocular forward-facing camera at various times and locations, including Suzhou and Urumqi, China, between October 2023 and January 2024. The dataset includes 1,086 images at various resolutions, which were resized to $1640\times 590$ for consistency.
Since foggy conditions are rare, we extended the dataset by incorporating foggy driving scenes from two YouTube videos(YouTube), resulting in an additional 337 images. The final FoggyLane dataset contains 1,423 images, covering six road scenes: Normal, Arrow, Crowd, Curve, Night, and Crossroad. The dataset was partitioned into training, validation and testing sets, with stratification to maintain a consistent distribution of scene types across both subsets.
Comparing FoggyLane with existing datasets, such as Caltech-Lanes(Caltech-Lanes) and Tusimple(Tusimple), which focus on single traffic scenarios, and LLAMAS(LLAMAS), which uses high-precision map annotations, we found that these datasets do not cover foggy weather scenarios. While ApolloScape (Apolloscape) and BDD100k (yu2020bdd100k) include multiple traffic scenarios, they lack instance-level annotations. Similarly, although CULane (pan2018spatial) and CurveLanes (xu2020curvelane) cover a wide range of traffic scenes, they do not include foggy conditions. Models trained on datasets limited to clear weather struggle in fog, where lane boundaries are blurred and image quality reduced. To fill this gap, FoggyLane offers a thorough solution for lane detection in foggy conditions with instance-level annotations.
Our dataset has two key advantages:
  • 1)
  • 2)
    Strict annotation standards: The dataset follows strict annotation guidelines, including marking the center of lanes, connecting lane markings at intersections, and prioritizing solid lines over grid lines, etc.(zhang2024robust). These rigorous annotations can significantly improve the accuracy and consistency of lane detection in foggy weather, thus enhancing the stability and safety of autonomous driving systems and providing a solid data foundation for the reliable deployment of autonomous technology.
4.1.2 FoggyCULane and FoggyTusimple
Currently, numerous datasets are available for lane detection. Among them, we select two widely adopted and representative benchmarks: CULane(pan2018spatial) and TuSimple(Tusimple). The CULane dataset (pan2018spatial) serves as an extensive and demanding benchmark, encompassing various scenarios such as highways, urban roads, and rural regions in Beijing. In contrast, the TuSimple dataset (Tusimple) primarily comprises daytime driving videos collected on American highways, featuring a variety of lighting conditions and complex traffic situations. Based on these datasets, we generate synthetic foggy versions, named FoggyCULane and FoggyTuSimple, to facilitate evaluation under adverse weather conditions.
The method for generating foggy images is inspired by the model (nie2022sensor), and the complete generation process is shown in Fig. 10. To be specific, the atmospheric scattering model can effectively utilize prior information to analyze the degradation mechanisms of image quality in foggy conditions. It models foggy images as a combination of scattered light and direct light based on optical principles, ultimately synthesizing a clear image into a foggy one. The mathematical modeling equation defined by Koschmieder (harald1924theorie) is represented as:
I(x)=J(x)t(x)+A(1t(x))I(x)=J(x)t(x)+A(1-t(x)) (13)
$t(x)=e^{-\beta d(x)}$ (14)
To calculate the corresponding $d(x)$, we use the S2R-DepthNet(chen2021S2R) that has good generalization performance across various applications. The S2R-DepthNet consists of three modules that decouple the structural and textural information of the input image. Then they suppress information irrelevant to depth and use structural information for depth prediction. The Structure Extraction Module (STE) breaks down the image to capture domain-independent structural details. The Depth-specific Attention (DSA) generates an attention map from the original image. Once processed through the DSA and STE Modules, as depicted in Fig. 10, the attention map and structural map are multiplied to produce a refined and all-encompassing depth-specific structure map. The Depth Prediction (DP) module is then responsible for estimating the depth.
According to the transmittance formula Eq. (14), we use the dark channel prior(he2011darkchannel) to calculate the average atmospheric light value $A$. In the non-sky area of an outdoor clear image, there is usually at least one very small or close to zero value in the RGB image for each pixel, which is called the dark channel of that pixel. Therefore, for any image $I$, the mathematical formula for its corresponding dark channel image $I_{\text{dark}}(x)$ can be defined as:
where $I_{\text{c}}(y)$ is the value of the c-th channel of image $I$ at pixel point $y$, and $\omega(x)$ is a small region around pixel point $x$. In the dark channel, find the top 0.1% brightest pixels and take the maximum value of their RGB channels as an approximation of the average atmospheric light value $A$. After obtaining the transmittance and the average atmospheric light value, the images of light fog, medium fog, and heavy fog can be obtained by fine-tuning the extinction coefficient $\beta$. The original images and the foggy images generated under different $\beta$ values are displayed in Figs. 11.
[…]
@@ 4.2 Evaluation Metric
Lane detection is typically approached as a semantic segmentation problem, assessed using common metrics. However, due to its distinct characteristics, lane detection also requires specialized evaluation metrics. Adhering to SCNN’s (pan2018spatial) framework, we link labeled or predicted lane points to form line segments and extend them into 30-pixel-wide regions. Subsequently, we compute the Intersection over Union (IoU), and if the IoU value exceeds a predefined threshold, typically set at 0.5, the prediction is considered correct. From this, we calculate Precision, Recall, and F1-score using the following formulas:
@@
Figure 12: Visualization results of CLRNet(zheng2022clrnet), FLAMNet(ran2023flamnet), PolarRCNN(tits4) and our method on FoggyLane testing set. The arrows indicate instances of lane lines that were not correctly detected by the corresponding method.
[…]
Figure 13: Visualization results of CLRNet(zheng2022clrnet), FLAMNet(ran2023flamnet), PolarRCNN(tits4) and our method on FoggyCULane testing set. The arrows indicate instances of lane lines that were not correctly detected by the corresponding method.
@@ 4.3 Implementation Details
To benchmark against existing methods(pan2018spatial; qin2020ultra; liu2021condlanenet; zheng2021resa; tabelini2021keep; liu2021end; zheng2022clrnet; ran2023flamnet; dai2024dbnet; tits4), we followed the training settings provided in the respective papers. For FoggyCULane and FoggyTusimple, we adopted the training parameters from the original CULane and Tusimple datasets, while for FoggyLane, we extended the training epochs to ensure reliable results
@@ 4.5 Ablation Study
We conducted ablation studies on the FoggyLane and FoggyCULane datasets to evaluate the effectiveness of the GFFM, DFFM, and PEM modules, using Swin-t as the backbone. As shown in Tables 5 and 6, each module improves the model’s performance. The GFFM module contributes the most, increasing the F1-score by 1.07 on FoggyLane and 0.41 on FoggyCULane, highlighting its ability to capture both local and global information. Adding the DFFM module further boosts the F1-score by 0.63 on FoggyLane and 0.26 on FoggyCULane. Finally, the PEM module improves edge detection, increasing the F1-score by 0.28 on FoggyLane and 0.14 on FoggyCULane.
@@
  Model GFFM DFFM PEM F1-score Gflops(G)
@@ 4.6 Model deployment
  Model GFFM DFFM PEM F1-score Gflops(G)
@@
Figure 15: Lane departure warning module within the active safety warning systems framework. The module detects unintentional lane departure using a forward-facing camera and edge computing system built with NVIDIA Jetson AGX Orin for real-time lane marking recognition. Upon detecting a departure trajectory (represented by yellow dotted lines), the system triggers multi-modal warnings including steering wheel vibration, seat vibration, and voice notification with corresponding warning icons on the dashboard monitor to proactively prevent accidents.
@@ 5 Conclusion
This study addresses the critical challenge of lane detection in foggy conditions, which is essential for improving road safety and enhancing Advanced Driver Assistance Systems (ADAS). To tackle the issue of insufficient publicly available datasets for foggy lane detection, we constructed the FoggyLane dataset, a real-world dataset specifically targeting lane detection in foggy environments, and synthesized two additional datasets, FoggyCULane and FoggyTusimple, to fill the gap in fog-specific data.
In response to the difficulties posed by foggy scenarios, we propose a robust Prior-Guided Dynamic Feature Fusion Transformer framework for real-time lane detection. This method incorporates a Global Feature Fusion Module to capture the relationship between local and global features in foggy images, a Dynamic Feature Fusion Module to model the structural and positional relationships of lane instances, and a Prior-Guided Edge Enhancement Module to recover lost edge details. Extensive experiments show that our approach significantly improves lane detection accuracy compared to other advanced methods, both in clear and foggy conditions, with F1-scores of 95.04%, 79.85%, and 96.95% on the FoggyLane, FoggyCULane, and FoggyTusimple datasets, respectively.
Moreover, with TensorRT acceleration, our method achieves an inference speed of 38.4 FPS on the NVIDIA Jetson AGX Orin, confirming its practical applicability in real-world scenarios. By improving the precision of lane detection, our framework contributes to active safety warning systems, helping prevent accidents in foggy conditions and enhancing road safety, particularly in challenging driving environments.